# local non-bossiness

### v4

@@ 5. Mechanisms under preferences over colleagues
Any stable mechanism $\Phi:{\mathcal{P}}\rightarrow{\mathcal{M}}$ induces a stable mechanism on $\mathcal{D}_{{\rm{c}}}$ by the rule that associates each $\unrhd\in\mathcal{D}_{{\rm{c}}}$ with the matching $\Phi(P(\unrhd))$. Evidently, many other stable mechanisms can be defined in $\mathcal{D}_{{\rm{c}}}$ using the information that students reveal about how they rank their colleagues. Despite this multiplicity, the uniqueness result of Alcalde and Barberà (1994) can be extended to the preference domain $\mathcal{D}_{{\rm{c}}}$.
@@
It follows from Theorem 3 that $\overline{\rm{DA}}$ is stable and strategy-proof on the preference domain $\mathcal{D}_{{\rm{c}}}$. By contradiction, assume that there is a stable and strategy-proof mechanism $\Omega:\mathcal{D}_{{\rm{c}}}\rightarrow{\mathcal{M}}$ such that $\Omega(\unrhd)\neq\overline{\rm{DA}}(\unrhd)$ for some $\unrhd\in\mathcal{D}_{{\rm{c}}}$. Since $\mathcal{D}_{{\rm{c}}}$ only includes school-lexicographic preferences, $\overline{\rm{DA}}(\unrhd)$ pairs each student to the most preferred alternative in $S\cup\{s_{0}\}$ that she can obtain in a stable outcome of $(S,N,\succ,q,\unrhd)$. [^15] That is, there is no stable matching $\mu$ such that $\mu(i)P_{i}(\unrhd)\overline{\rm{DA}}_{i}(\unrhd)$ for some $i\in N$. Thus, $\overline{\rm{DA}}_{i}(\unrhd)\rhd_{i}\Omega_{i}(\unrhd)$ for some $i\in N$. In particular, $\overline{\rm{DA}}_{i}(\unrhd)$ is a school.
Let $P^{\prime}_{i}$ be the preferences defined on $S\cup\{s_{0}\}$ for which $\overline{\rm{DA}}_{i}(\unrhd)$ is the only acceptable school (i.e., $s_{0}P^{\prime}_{i}s$ for all school $s\neq\overline{\rm{DA}}_{i}(\unrhd)$). Fix $\unrhd^{\prime}=(\unrhd^{\prime}_{i},\unrhd_{-i})\in\mathcal{D}_{{\rm{c}}}$ such that $P_{i}(\unrhd^{\prime})=P_{i}^{\prime}$. Since the problems $(S,N,\succ,q,\unrhd^{\prime})$ and $(S,N,\succ,q,(P^{\prime}_{i},P_{-i}(\unrhd)))$ have the same stable matchings, and $\overline{\rm{DA}}(\unrhd)$ is stable under $(P^{\prime}_{i},P_{-i}(\unrhd))$, the definition of $P^{\prime}_{i}$ and the Rural Hospital Theorem (Roth, 1986) imply that $\Omega_{i}(\unrhd^{\prime}_{i},\unrhd_{-i})=\overline{\rm{DA}}_{i}(\unrhd)$. Therefore, $\Omega_{i}(\unrhd^{\prime}_{i},\unrhd_{-i})\rhd_{i}\Omega_{i}(\unrhd)$, which contradicts the strategy-proofness of $\Omega$. ∎
@@
Let ${\mathcal{D}}_{1}$ be the preference domain in which the student $1$ has school-lexicographic preferences and the other students have school-lexicographic preferences over colleagues. We claim that no stable mechanism defined in ${\mathcal{D}}_{1}$ is strategy-proof.
@@
$P_{1}(\unrhd):s_{3}\ldots;\quad\quad\quad P_{2}(\unrhd):s_{2},s_{1},\ldots;\quad\quad\quad P_{3}(\unrhd):s_{1},s_{2},\ldots;$
$P_{4}(\unrhd):s_{1},\ldots;\quad\quad\quad P_{5}(\unrhd):s_{4},\ldots.$
Notice that $[N,S,\succ,q,P(\unrhd)]$ has only two stable matchings:
@@
Since there are no restrictions on how the student $1$ ranks $\mu$ and $\eta$, because she is assigned to the same school in both matchings and has school-lexicographic preferences, assume that she strictly prefers $\mu$ to $\eta$ under $\unrhd_{1}$.
It is not difficult to verify that $\mu$ and $\eta$ are the only stable matchings of $[N,S,\succ,q,\unrhd]$. Hence, if a mechanism $\Phi:\mathcal{D}_{1}\rightarrow{\mathcal{M}}$ is stable, then $\Phi(\unrhd)\in\{\mu,\eta\}$.
Suppose that $\Phi(\unrhd)=\mu$. If $\unrhd^{\prime}_{2}$ is a school-lexicographic preference over colleagues such that $P_{2}^{\prime}\equiv P_{2}(\unrhd^{\prime}_{2},\unrhd_{-2})$ satisfies $P^{\prime}_{2}:s_{2},s_{4},\ldots$, then $\eta$ is the only stable matching of $[N,S,\succ,q,(\unrhd^{\prime}_{2},\unrhd_{-2})]$. Hence, the student $2$ has incentives to manipulate $\Phi$, because $s_{2}P_{2}(\unrhd)s_{1}$ implies that $\Phi(\unrhd^{\prime}_{2},\unrhd_{-2})=\eta\rhd_{2}\mu=\Phi(\unrhd)$.
Suppose that $\Phi(\unrhd)=\eta$. If $\unrhd^{\prime}_{1}$ is such that $P_{1}^{\prime}\equiv P_{1}(\unrhd^{\prime}_{1},\unrhd_{-1})$ satisfies $P^{\prime}_{1}:s_{1},s_{3},\ldots$, then $\mu$ is the only stable matching of $[N,S,\succ,q,(\unrhd^{\prime}_{1},\unrhd_{-1})]$. Hence, the student $1$ has incentives to manipulate $\Phi$, because $\Phi(\unrhd^{\prime}_{1},\unrhd_{-1})=\mu\rhd_{1}\eta=\Phi(\unrhd).$
@@
Given a school choice context $[N,S,\succ,q]$ and a domain ${\mathcal{D}}^{\prime}$ such that ${\mathcal{D}}_{\rm{c}}\subsetneq{\mathcal{D}}^{\prime}\subseteq{\mathcal{D}}$, to ensure that no mechanism defined in ${\mathcal{D}}^{\prime}$ is stable and strategy-proof it is crucial that $\succ$ has an Ergin-cycle. [^16] The priority profile of 2 has an Ergin-cycle characterized by $(s,s^{\prime})=(s_{1},s_{2})$, $(i,j,k)=(2,1,3)$, $N_{s}=\{4\}$, and $N_{s^{\prime}}=\emptyset$. Indeed, the mechanism $\overline{\rm{DA}}$ is stable and strategy-proof in ${\mathcal{D}}$ whenever $\succ$ is acyclic (see Duque and Torres-Martínez, 2023).
@@ 6. Concluding remarks
In this paper, we have shown that the bossiness of the ${\rm{DA}}$ mechanism is limited. When ${\rm{DA}}$ is implemented, it is well known that a student can affect another student’s assignment by changing her preferences without modifying her school. We have demonstrated that this holds only for students not assigned to the same school. In other words, under ${\rm{DA}}$, a student cannot change her schoolmates without changing her own school. Additionally, for any strategy-proof mechanism, local non-bossiness guarantees that no coalition of students assigned to the same school can misrepresent their preferences to either improve their assignments or maintain their school while modifying their colleagues Since ${\rm{DA}}$ is not group strategy-proof, our result implies that a successful manipulating coalition of DA must include students from different schools.
Local non-bossiness is not only interesting in itself but also because of its application to school choice problems with externalities. In this framework, even when students prioritize their assigned school first and then consider the assignment of others, a stable and strategy-proof mechanism may not exist (Duque and Torres-Martínez, 2023). However, if we restrict preferences to be such that only the school and the schoolmates matter, it turns out that ${\rm{DA}}$ induces the only stable and strategy-proof mechanism. The crucial property behind this result is precisely its local non-bossiness.
Although we have focused on many-to-one matching problems with responsive preferences, it might be interesting to investigate the role of local non-bossiness in more general frameworks, as many-to-one matching models with contracts (Hatfield and Milgrom, 2005) or many-to-many matching problems (Echenique and Oviedo, 2006). These extensions are of interest due to their potential applications to more complex labor or educational markets, such as school choice with affirmative action or part-time work assignments (e.g., medical interns in hospitals or teachers in public schools). However, to guarantee that a mechanism $\Phi$ is locally non-bossy using our approach, the following three properties are crucial: (i) $\Phi$ must be a stable mechanism; (ii) $\Phi$ must assign all students to the most/least preferred school they can attend in a stable matching; and (iii) any school that does not fill its quota in some stable matching must enroll the same set of students in any other stable outcome. Therefore, any attempt to extend our results to other contexts will likely require restricting the preference domain to ensure that properties analogous to (i)-(iii) hold. This is a matter for future research.
Appendix A local non-bosiness and its implications
@@
Given $P\in{\mathcal{P}}$ and $s\in S\cup\{s_{0}\}$, suppose that there is a coalition $C=\{i_{1},\ldots,i_{r}\}$ contained in $\Phi_{s}(P)$ such that $\Phi_{i}(P^{\prime}_{C},P_{-C})R_{i}\,\Phi_{i}(P)$ for all $i\in C$ and for some $P^{\prime}_{C}=(P^{\prime}_{j})_{j\in C}$. We will prove that $\Phi_{i}(P^{\prime}_{C},P_{-C})=\,\Phi_{i}(P)$ for all $i\in C$.
For each $i\in\Phi_{s}(P)$, let $\widehat{P}_{i}$ be the preference relation that places $\Phi_{i}(P^{\prime}_{C},P_{-C})$ at the top and keeps the other schools in the order induced by $P_{i}$. Suppose that $\Phi_{i_{1}}(P)\neq\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})$. The strategy-proofness of $\Phi$ implies that $\Phi_{i_{1}}(P)P_{i_{1}}\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})$ and $\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})\widehat{P}_{i_{1}}\Phi_{i_{1}}(P).$ If $\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})\neq\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})$, the definition of $\widehat{P}_{i_{1}}$ implies that $\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})P_{i}\Phi_{i_{1}}(P)$. Hence, $\Phi_{i_{1}}(P)P_{i_{1}}\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})P_{i_{1}}\Phi_{i_{1}}(P)$, which is not possible. As a consequence, $\Phi_{i_{1}}(P)\neq\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})$ ensures that $\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})=\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})$, which in turn implies that $\Phi_{i_{1}}(P)P_{i_{1}}\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})$, a contradiction. We conclude that $\Phi_{i_{1}}(P)=\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})$, and the local non-bossiness of $\Phi$ implies that $\Phi_{i}(P)=\Phi_{i}(\widehat{P}_{i_{1}},P_{-i_{1}})$ for all $i\in\Phi_{s}(P)$. Repeating the argument for students $i_{2}$,…, $i_{r}$ we obtain that $\Phi_{i}(P)=\Phi_{i}(\widehat{P}_{C},P_{-C})$ for all $i\in\Phi_{s}(P).$
Since $\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})$ is the most preferred alternative under $\widehat{P}_{i_{1}}$, it follows that $\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})\widehat{R}_{i_{1}}\Phi_{i_{1}}(\widehat{P}_{i_{1}},P^{\prime}_{C\setminus\{i_{1}\}},P_{-C})$. If $\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})\neq\Phi_{i_{1}}(\widehat{P}_{i_{1}},P^{\prime}_{C\setminus\{i_{1}\}},P_{-C})$, then $\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})\widehat{P}_{i_{1}}\Phi_{i_{1}}(\widehat{P}_{i_{1}},P^{\prime}_{C\setminus\{i_{1}\}},P_{-C})$, which contradicts the strategy-proofness of $\Phi$. Moreover, as $\Phi$ is locally non-bossy, $\Phi_{i}(P^{\prime}_{C},P_{-C})=\Phi_{i}(\widehat{P}_{i_{1}},P^{\prime}_{C\setminus\{i_{1}\}},P_{-C})$ for all $i\in\Phi_{s}(P)$. Repeating the argument for $i_{2}$,…, $i_{r}$, we obtain that $\Phi_{i}(P^{\prime}_{C},P_{-C})=\Phi_{i}(\widehat{P}_{C},P_{-C})$ for all $i\in\Phi_{s}(P).$ Therefore, $\Phi_{i}(P^{\prime}_{C},P_{-C})=\Phi_{i}(P)$ for every student $i\in C$. ∎
@@
Afacan (2012) introduces the following extension of non-bossiness to group of students: a mechanism $\Phi$ is group non-bossy when for any $P\in{\mathcal{P}}$ and $C\subseteq N$, if there exists $P^{\prime}_{C}\in{\mathcal{L}}^{|C|}$ such that $\Phi_{i}(P)=\Phi_{i}(P^{\prime}_{C},P_{-C})$ for all $i\in C$, then $\Phi(P)=\Phi(P^{\prime}_{C},P_{-C})$.
We will consider a local version of this property: a mechanism $\Phi$ is locally group non-bossy when for any $s\in S\cup\{s_{0}\}$, $P\in{\mathcal{P}}$, and $C\subseteq\Phi_{s}(P)$, if there exists $P^{\prime}_{C}\in{\mathcal{L}}^{|C|}$ such that $\Phi_{i}(P)=\Phi_{i}(P^{\prime}_{C},P_{-C})$ for all $i\in C$, then $\Phi_{s}(P)=\Phi_{s}(P^{\prime}_{C},P_{-C})$.
Hence, local group non-bossiness ensures that no coalition of students assigned to the same school can change the assignment of some of their colleagues by reporting different preferences, without changing their school.
@@
The converse of Lemma 2 does not hold. Indeed, consider the Boston mechanism, also known as the Immediate Acceptance mechanism. This mechanism runs similarly to DA, with the difference that at each step accepted students are definitively matched to the school. Although it is not strategy-proof (Abdulkadiroğlu and Sönmez, 2003), it is locally group non-bossy. Formally, denoting by ${\mathcal{B}}:{\mathcal{P}}\rightarrow{\mathcal{M}}$ this mechanism, suppose that ${\mathcal{B}}_{i}(P)={\mathcal{B}}_{i}(P^{\prime}_{C},P_{-C})=s$ for all student in a coalition $C\subseteq{\mathcal{B}}_{s}(P)$. Let $\widehat{P}_{C}=(\widehat{P}_{i})_{i\in C}$ be such that $s$ is the most preferred alternative under each $\widehat{P}_{i}$. It is not difficult to verify that ${\mathcal{B}}_{s}(P)={\mathcal{B}}_{s}(\widehat{P}_{C},P_{-C})={\mathcal{B}}_{s}(P^{\prime}_{C},P_{-C})$.
@@
Since the school-optimal stable mechanism in $[N,S,q,\succ,\overline{P}]$ is
@@
Since the Gale’s Top Trading Cycles mechanism (Shapley and Scarf, 1974) is non-bossy and unstable, it follows from 3 that local non-bossiness and stability are independent properties.
Bando and Imamura (2016) introduce the following property: a mechanism $\Phi$ is weak non-bossy when for any $P\in{\mathcal{P}}$, $i\in N$, and $P^{\prime}_{i}\in{\mathcal{L}}$, $\Phi_{i}(P)=\Phi_{i}(P^{\prime}_{i},P_{-i})$ implies that $\Phi_{s_{0}}(P)=\Phi_{s_{0}}(P^{\prime}_{i},P_{-i})$. Hence, when a mechanism is weak non-bossy, no student can modify the set of unassigned students without changing her own school.
Since DA satisfies weak non-bossiness (Bando and Imamura, 2016; Theorem 3), it is interesting to compare it with local non-bossiness.
@@
Weak non-bossiness is independent of local non-bossiness and strategy-proofness.
@@
On the one hand, it is not difficult to verify that the mechanism described in 3 is weak non-bossy. However, this mechanism is neither locally non-bossy nor strategy-proof. On the other hand, if $N=\{1,2,3\}$, $S=\{s_{1},s_{2}\}$, and $(q_{s_{1}},q_{s_{2}})=(2,1),$ regardless of the schools’ priority profile $(\succ_{s_{1}},\succ_{s_{2}})$, the mechanism $\Omega:{\mathcal{P}}\rightarrow{\mathcal{M}}$ that is defined by
$\Omega(P)=\left\{\begin{array}[]{ll}((1,s_{1}),(2,s_{1}),(3,s_{2})),&\quad\quad\mbox{when}\,\,s_{2}P_{1}s_{1},\\ ((1,s_{1}),(2,s_{1}),(3,s_{0})),&\quad\quad\mbox{in another case},\end{array}\right.$

### v5

@@ 5. Mechanisms under preferences over colleagues
Every stable mechanism $\Phi:{\mathcal{P}}\rightarrow{\mathcal{M}}$ induces a stable mechanism on $\mathcal{D}_{{\rm{c}}}$ by the rule that associates each $\unrhd\in\mathcal{D}_{{\rm{c}}}$ with the matching $\Phi(P(\unrhd))$. Evidently, many other stable mechanisms can be defined in $\mathcal{D}_{{\rm{c}}}$ using the information that students reveal about how they rank their colleagues. Despite this multiplicity, the uniqueness result of Alcalde and Barberà (1994) can be extended to the preference domain $\mathcal{D}_{{\rm{c}}}$.
@@
It follows from Theorem 3 that $\overline{\rm{DA}}$ is stable and strategy-proof on the preference domain $\mathcal{D}_{{\rm{c}}}$. By contradiction, assume that there is a stable and strategy-proof mechanism $\Omega:\mathcal{D}_{{\rm{c}}}\rightarrow{\mathcal{M}}$ such that $\Omega(\unrhd)\neq\overline{\rm{DA}}(\unrhd)$ for some $\unrhd\in\mathcal{D}_{{\rm{c}}}$. Since $\mathcal{D}_{{\rm{c}}}$ only includes school-lexicographic preferences, $\overline{\rm{DA}}(\unrhd)$ pairs each student to the most preferred alternative in $S\cup\{s_{0}\}$ that she can obtain in a stable outcome of $(S,N,\succ,q,\unrhd)$. [^17] That is, there is no stable matching $\mu$ such that $\mu(i)\,P_{i}(\unrhd_{i})\,\overline{\rm{DA}}_{i}(\unrhd)$ for some $i\in N$. Thus, $\overline{\rm{DA}}_{i}(\unrhd)\rhd_{i}\Omega_{i}(\unrhd)$ for some $i\in N$. In particular, $\overline{\rm{DA}}_{i}(\unrhd)$ is a school.
Let $P^{\prime}_{i}$ be the preferences defined on $S\cup\{s_{0}\}$ for which $\overline{\rm{DA}}_{i}(\unrhd)$ is the only acceptable school (i.e., $s_{0}P^{\prime}_{i}s$ for all school $s\neq\overline{\rm{DA}}_{i}(\unrhd)$). Fix $\unrhd^{\prime}=(\unrhd^{\prime}_{i},\unrhd_{-i})\in\mathcal{D}_{{\rm{c}}}$ such that $P_{i}(\unrhd^{\prime}_{i})=P_{i}^{\prime}$. Since the problems $(S,N,\succ,q,\unrhd^{\prime})$ and $(S,N,\succ,q,(P^{\prime}_{i},P_{-i}(\unrhd)))$ have the same stable matchings, and $\overline{\rm{DA}}(\unrhd)$ is stable under $(P^{\prime}_{i},P_{-i}(\unrhd))$, the definition of $P^{\prime}_{i}$ and the Rural Hospital Theorem (Roth, 1984b; Gale and Sotomayor, 1985) imply that $\Omega_{i}(\unrhd^{\prime}_{i},\unrhd_{-i})=\overline{\rm{DA}}_{i}(\unrhd)$. Therefore, $\Omega_{i}(\unrhd^{\prime}_{i},\unrhd_{-i})\rhd_{i}\Omega_{i}(\unrhd)$, which contradicts the strategy-proofness of $\Omega$. ∎
@@
Let ${\mathcal{D}}_{1}$ be the preference domain in which student $1$ has school-lexicographic preferences and the other students have school-lexicographic preferences over colleagues. We claim that no stable mechanism defined in ${\mathcal{D}}_{1}$ is strategy-proof.
@@
$P_{1}(\unrhd_{1}):\,s_{3}\ldots;\quad\quad\quad P_{2}(\unrhd_{2}):\,s_{2},s_{1},\ldots;\quad\quad\quad P_{3}(\unrhd_{3}):\,s_{1},s_{2},\ldots;$
$P_{4}(\unrhd_{4}):\,s_{1},\ldots;\quad\quad\quad P_{5}(\unrhd_{5}):\,s_{4},\ldots.$
Note that $N,S,\succ,q,P(\unrhd)]$ has only two stable matchings:
@@
Since there are no restrictions on how student $1$ ranks $\mu$ and $\eta$, because she is assigned to the same school in both matchings and has school-lexicographic preferences, assume that she strictly prefers $\mu$ to $\eta$ under $\unrhd_{1}$.
It is not difficult to verify that $\mu$ and $\eta$ are the only stable matchings of $N,S,\succ,q,\unrhd]$. Hence, if a mechanism $\Phi:\mathcal{D}_{1}\rightarrow{\mathcal{M}}$ is stable, then $\Phi(\unrhd)\in\{\mu,\eta\}$.
Suppose that $\Phi(\unrhd)=\mu$. If $\unrhd^{\prime}_{2}$ is a school-lexicographic preference over colleagues such that $P_{2}^{\prime}\equiv P_{2}(\unrhd^{\prime}_{2})$ satisfies $P^{\prime}_{2}:s_{2},s_{4},\ldots$, then $\eta$ is the only stable matching of $N,S,\succ,q,(\unrhd^{\prime}_{2},\unrhd_{-2})]$. Hence, the student $2$ has incentives to manipulate $\Phi$, because $s_{2}P_{2}(\unrhd_{2})s_{1}$ implies that $\Phi(\unrhd^{\prime}_{2},\unrhd_{-2})=\eta\rhd_{2}\mu=\Phi(\unrhd)$.
Suppose that $\Phi(\unrhd)=\eta$. If $\unrhd^{\prime}_{1}$ is such that $P_{1}^{\prime}\equiv P_{1}(\unrhd^{\prime}_{1})$ satisfies $P^{\prime}_{1}:s_{1},s_{3},\ldots$, then $\mu$ is the only stable matching of $N,S,\succ,q,(\unrhd^{\prime}_{1},\unrhd_{-1})]$. Hence, student $1$ has incentives to manipulate $\Phi$, because $\Phi(\unrhd^{\prime}_{1},\unrhd_{-1})=\mu\rhd_{1}\eta=\Phi(\unrhd).$
@@
Given a school choice context $N,S,\succ,q]$ and a domain $\mathcal{D}$$^{\prime}$ such that ${\mathcal{D}}_{\rm{c}}\subsetneq{\mathcal{D}}^{\prime}\subseteq{\mathcal{D}}$, to ensure that no mechanism defined in $\mathcal{D}$$^{\prime}$ is stable and strategy-proof it is crucial that $\succ$ has a cycle in the sense of Ergin (2002). [^18] See the previous section for the definition of Ergin-cycle. The priority profile of 2 has an Ergin-cycle characterized by $(s,s^{\prime})=(s_{1},s_{2})$, $(i,j,k)=(2,1,3)$, $N_{s}=\{4\}$, and $N_{s^{\prime}}=\emptyset$. Indeed, the mechanism $\overline{\rm{DA}}$ is stable and strategy-proof in $\mathcal{D}$} whenever $\succ$ is Ergin-acyclic (see Duque and Torres-Martínez, 2023).
@@ 6. Concluding remarks
In this paper, we have shown that the bossiness of the student-optimal stable mechanism is limited. When $\rm{DA}$} is implemented, it is well known that a student can affect another student’s assignment by changing her preferences without modifying her school. We have demonstrated that this holds only for students who are not assigned to the same school. In other words, under $\rm{DA}$}, a student cannot change her schoolmates without changing her own school. Additionally, for every strategy-proof mechanism, local non-bossiness guarantees that no coalition of students assigned to the same school can misrepresent their preferences to either improve their assignments or maintain their school while modifying their colleagues. Since $\rm{DA}$} is not group strategy-proof, our result implies that every successful manipulating coalition of DA includes students from different schools.
Local non-bossiness is not only interesting in itself but also because of its application to school choice problems with externalities. In this framework, even when students prioritize their assigned school first and then consider the assignment of others, a stable and strategy-proof mechanism may not exist (Duque and Torres-Martínez, 2023). However, if we restrict preferences to be such that only the own school and the schoolmates matter, it turns out that $\rm{DA}$} induces the only stable and strategy-proof mechanism. The crucial property behind this result is precisely its local non-bossiness.
Appendix A local non-bossiness and its implications
We analyze the relationship between local non-bossiness and other incentive properties, such as strategy-proofness, (local) group strategy-proofness, weak non-bossiness (Bando and Imamura, 2016), and a local notion of group non-bossiness (Afacan, 2012). The following diagram summarizes the causal relationships between these incentive properties that we formalize in this appendix.
[…]
Figure 1. On local non-bossiness and other incentive properties.
@@
Given $P\in{\mathcal{P}}$ and $s\in S\cup\{s_{0}\}$, suppose that there is a coalition $C=\{i_{1},\ldots,i_{r}\}$ contained in $\Phi_{s}(P)$ such that $\Phi_{i}(P^{\prime}_{C},P_{-C})R_{i}\,\Phi_{i}(P)$ for all $i\in C$ and for some $P^{\prime}_{C}=(P^{\prime}_{j})_{j\in C}$. We will prove that $\Phi_{i}(P^{\prime}_{C},P_{-C})=\,\Phi_{i}(P)$ for all $i\in C$. For each $i\in\Phi_{s}(P)$, let $\widehat{P}_{i}$ be the preference relation that places $\Phi_{i}(P^{\prime}_{C},P_{-C})$ at the top and keeps the other schools in the order induced by $P_{i}$.
Suppose that $\Phi_{i_{1}}(P)\neq\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})$. The strategy-proofness of $\Phi$ implies that $\Phi_{i_{1}}(P)P_{i_{1}}\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})$ and $\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})\widehat{P}_{i_{1}}\Phi_{i_{1}}(P).$ If $\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})\neq\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})$, the definition of $\widehat{P}_{i_{1}}$ implies that $\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})P_{i}\Phi_{i_{1}}(P)$. Hence, $\Phi_{i_{1}}(P)P_{i_{1}}\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})P_{i_{1}}\Phi_{i_{1}}(P)$, which is not possible. As a consequence, $\Phi_{i_{1}}(P)\neq\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})$ ensures that $\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})=\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})$, which in turn implies that $\Phi_{i_{1}}(P)P_{i_{1}}\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})$, a contradiction.
Since $\Phi_{i_{1}}(P)=\Phi_{i_{1}}({\widehat{P}}_{i_{1}},P_{-i_{1}})$, the local non-bossiness of $\Phi$ implies that $\Phi_{i}(P)=\Phi_{i}(\widehat{P}_{i_{1}},P_{-i_{1}})$ for all $i\in\Phi_{s}(P)$. Repeating the arguments above for students $i_{2}$,…, $i_{r}$ we obtain that $\Phi_{i}(P)=\Phi_{i}(\widehat{P}_{C},P_{-C})$ for all $i\in\Phi_{s}(P).$
On the other hand, as $\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})$ is the most preferred alternative under $\widehat{P}_{i_{1}}$, it follows that $\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})\widehat{R}_{i_{1}}\Phi_{i_{1}}(\widehat{P}_{i_{1}},P^{\prime}_{C\setminus\{i_{1}\}},P_{-C})$. If $\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})\neq\Phi_{i_{1}}(\widehat{P}_{i_{1}},P^{\prime}_{C\setminus\{i_{1}\}},P_{-C})$, then $\Phi_{i_{1}}(P^{\prime}_{C},P_{-C})\widehat{P}_{i_{1}}\Phi_{i_{1}}(\widehat{P}_{i_{1}},P^{\prime}_{C\setminus\{i_{1}\}},P_{-C})$, which contradicts the strategy-proofness of $\Phi$. Moreover, as $\Phi$ is locally non-bossy, $\Phi_{i}(P^{\prime}_{C},P_{-C})=\Phi_{i}(\widehat{P}_{i_{1}},P^{\prime}_{C\setminus\{i_{1}\}},P_{-C})$ for all $i\in\Phi_{s}(P)$. Repeating this argument for students $i_{2}$,…, $i_{r}$, we obtain that $\Phi_{i}(P^{\prime}_{C},P_{-C})=\Phi_{i}(\widehat{P}_{C},P_{-C})$ for all $i\in\Phi_{s}(P).$
Therefore, $\Phi_{i}(P^{\prime}_{C},P_{-C})=\Phi_{i}(P)$ for every student $i\in C$. ∎
@@
Afacan (2012) introduces the following extension of non-bossiness to groups of students: a mechanism $\Phi$ is group non-bossy when for every $P\in{\mathcal{P}}$ and $C\subseteq N$, if there exists $P^{\prime}_{C}\in{\mathcal{L}}^{|C|}$ such that $\Phi_{i}(P)=\Phi_{i}(P^{\prime}_{C},P_{-C})$ for all $i\in C$, then $\Phi(P)=\Phi(P^{\prime}_{C},P_{-C})$.
We will consider a local version of this property: a mechanism $\Phi$ is locally group non-bossy when for every $s\in S\cup\{s_{0}\}$, $P\in{\mathcal{P}}$, and $C\subseteq\Phi_{s}(P)$, if there exists $P^{\prime}_{C}\in{\mathcal{L}}^{|C|}$ such that $\Phi_{i}(P)=\Phi_{i}(P^{\prime}_{C},P_{-C})$ for all $i\in C$, then $\Phi_{s}(P)=\Phi_{s}(P^{\prime}_{C},P_{-C})$. Hence, local group non-bossiness ensures that no coalition of students assigned to the same school can change the assignment of some of their colleagues by reporting different preferences and without changing their school.
@@
The converse of Lemma 2 does not hold. Indeed, consider the Boston mechanism, also known as the Immediate Acceptance mechanism. This mechanism runs similarly to DA, with the difference that at each step accepted students are definitively matched to schools. Although it is not strategy-proof (Abdulkadiroğlu and Sönmez, 2003), it is locally group non-bossy. Formally, denoting by ${\mathcal{B}}:{\mathcal{P}}\rightarrow{\mathcal{M}}$ this mechanism, suppose that ${\mathcal{B}}_{i}(P)={\mathcal{B}}_{i}(P^{\prime}_{C},P_{-C})=s$ for all students in a coalition $C\subseteq{\mathcal{B}}_{s}(P)$. Let $\widehat{P}_{C}=(\widehat{P}_{i})_{i\in C}$ be such that $s$ is the most preferred alternative under each $\widehat{P}_{i}$. It is not difficult to verify that ${\mathcal{B}}_{s}(P)={\mathcal{B}}_{s}(\widehat{P}_{C},P_{-C})={\mathcal{B}}_{s}(P^{\prime}_{C},P_{-C})$.
@@
Since the school-optimal stable matching of $N,S,\succ,q,\overline{P}]$ is
@@
Since Gale’s Top Trading Cycles mechanism (Shapley and Scarf, 1974) is non-bossy and unstable, 3 implies that local non-bossiness and stability are independent properties.
Bando and Imamura (2016) introduce the following property: a mechanism $\Phi$ is weakly non-bossy when for every $P\in{\mathcal{P}}$, $i\in N$, and $P^{\prime}_{i}\in{\mathcal{L}}$, $\Phi_{i}(P)=\Phi_{i}(P^{\prime}_{i},P_{-i})$ implies that $\Phi_{s_{0}}(P)=\Phi_{s_{0}}(P^{\prime}_{i},P_{-i})$. Hence, when a mechanism is weakly non-bossy, no student can modify the set of unassigned students without changing her own school. Since DA satisfies weak non-bossiness (Bando and Imamura, 2016; Theorem 3), it is interesting to compare it with local non-bossiness.
@@
Weak non-bossiness is independent of both local non-bossiness and strategy-proofness.
@@
On the one hand, it is not difficult to verify that the mechanism described in 3 is weakly non-bossy. However, this mechanism is neither locally non-bossy nor strategy-proof. On the other hand, if $N=\{1,2,3\}$, $S=\{s_{1},s_{2}\}$, and $(q_{s_{1}},q_{s_{2}})=(2,1),$ regardless of the schools’ priority profile $(\succ_{s_{1}},\succ_{s_{2}})$, the mechanism $\Omega:{\mathcal{P}}\rightarrow{\mathcal{M}}$ defined by
$\Omega(P)=\left\{\begin{array}[]{ll}((1,s_{1}),(2,s_{1}),(3,s_{2})),&\quad\quad\mbox{when}\,\,s_{2}P_{1}s_{1},\\ ((1,s_{1}),(2,s_{1}),(3,s_{0})),&\quad\quad\mbox{otherwise},\end{array}\right$.